Lannister Manufacturing has a target debt-equity ratio of 0.66. Its cost of equity is 20 percent, and its cost of debt is 12 percent. If the tax rate is 33 percent, what is the company's WACC? Multiple Choice

16.01%

14.48%

12.8%

15.24%
Answer #1

After-tax cost of debt=12*(1-tax rate)

=12*(1-0.33)=8.04%

Debt-equity ratio=Debt/equity

Hence debt=0.66equity

Let equity be $x

Debt=$0.66x

Total=$1.66x

WACC=Respective costs*Respective weight

=(x/1.66x*20)+(0.66x/1.66x*8.04)

=15.24%(Approx).